• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kuchina

    Log entry time 08:00:19 on November 5,2006

    Entry number 188912

    keyword=Owl shift summary

    Troublesome shift:

    12am No beam goes to unstable beam at the beginning of the shift.
    MCC could not deliver more than 45uA, beam would trip every time they go beyond 45uA.
    1:00am Beam is stabilized.
    1:01am Coda problem, had to restart coda. Didn't help.
    Rebooted ROCs. Restarted coda. Started run (#1969).
    2:05am Beam is off for about an hour.
    3:06am Beam comes back,
    3:10am IOC16 dies, rebooted. Right Dipole NMR rebooted.
    3:14am MCC called: Vertical raster HV has been failed.
    Spent some time trying to get the help from tech on call.
    4:30am With no help from tech and no info on how to fix a problem, started data taking at low (5uA) current.